OpenStudy (ledah):

Simplify: (5xy)(2x^2)(y^3) 7x^3y^4 8x^2y^3 10x^3y^4 none of the above

Parth (parthkohli):

\( \color{Black}{\Rightarrow 5 \times x \times y \times 2 \times x \times x \times y \times y \times y }\)

Parth (parthkohli):

You just group 'em

Parth (parthkohli):

You can multiply 5,2... the y's, the x's :)

OpenStudy (ledah):

group the exponents?

Parth (parthkohli):

Yeah

Parth (parthkohli):

How many times does x repeat? that'd be the exponent of x. Similar for y :)

OpenStudy (ledah):

so its c right?

Parth (parthkohli):

Yes. c is right :)

OpenStudy (ledah):

I'm just checking my answer for this one..

OpenStudy (ledah):

okay thank you.

OpenStudy (anonymous):

\[(5xy)(2x^2)(y^3)\] Remember you can add like exponentials that are being multiplied \[10x^3y^4\] Rule (sorry the formatting isn't showing correctly): \[(\eta^A)*(\eta^B)=\eta^(A+B)\]